Log 1/8cosec 2 π/8 sin 2 3π/8 =?


Log1/8cosec2 π/8  sin2 3π/8 =?


Grade:12th Pass

1 Answers

APURV GOEL
12 years ago

Log1/8cosec2 π/8  sin2 3π/8 = Log 1/(2√2 sinΠ/8) sin3Π/8

                                                = Log 1/(2√2 sinΠ/8) cosΠ/8

we have sinΠ/4 = 2.sinΠ/8.cosΠ/8

so sinΠ/8 = (sinΠ/4)/2.cosΠ/8

so 1/(2√2 sinΠ/8) = cosΠ/8

hence Log 1/(2√2 sinΠ/8) cosΠ/8 = LogcosΠ/8cosΠ/8 = 1